.can you please help me solve this problem. i am stuck.
Find the length of a rectangular lot with a perimeter of 96 meters if the
length is 8 meters more than the width. (P = 2L + 2W)

Found 2 solutions by Alan3354, rfer:
Answer by Alan3354(69443)

Find the length of a rectangular lot with a perimeter of 96 meters if the
length is 8 meters more than the width. (P = 2L + 2W)
-----------
L + W = 48
L - W = 8
---------- Add
Can you finish it?

Answer by rfer(16322)

2(w+8)+2w=96
2w+16+2w=96
4w=96-16
4w=80
w=80/4
w=20
w+8=28